Let {an}∞n=1 be a non-decreasing (resp. non-increasing) sequence which converges to a. Then prove that an ≤ a (resp. a ≤ an) for every n ∈ N


1
Expert's answer
2021-02-01T06:37:12-0500

Let "(a_n)_{n=1}^{\infty}" be a non-decreasing (resp. non-increasing) sequence which converges to

"a", so we have that

"\lim \limits_{n \to \infty} a_n = a" "... (*)"

Also, "a_n \leq a_{n+1}" (resp. "a_n \geq a_{n+1}) \forall n \geq 1"

So we have from "(*)" that "\lim \limits_{n \to \infty} a_n \leq a"

(resp. "\lim \limits_{n \to \infty} a_n \geq a" ).

And since "(a_n)_{n=1}^{\infty}" is non-decreasing (resp. non-increasing), we have that "a_n \leq a" (resp. "a_n \geq a" ) As desired